# “Confidently Nonsensical?”: A Critical Survey on the Perspectives and Challenges of ‘Hallucinations’ in NLP

### 11 Apr 2024 v1

@@ Abstract
We investigate how hallucination in large language models (LLM) is characterized in peer-reviewed literature, using a critical examination of 103 publications across NLP research. Through a comprehensive review of sociological and technological literature, we identify a lack of agreement with the term ‘hallucination’. Additionally, we conduct a survey with 171 practitioners from the field of NLP and AI to capture varying perspectives on hallucination. Our analysis underscores the necessity for explicit definitions and frameworks outlining hallucination within NLP, highlighting potential challenges, and our survey inputs provide a thematic understanding of the influence and ramifications of hallucination in society.
@@ 1 Introduction
The utilization of Natural Language Processing (NLP), particularly in language generation, has witnessed significant advancement in recent months, owing to the emergence of highly proficient large language multi-modal models like GPT-4 and Gemini Gautam et al. (2024). These models have evolved beyond NLP tools used in technical domains to become sociotechnical systems, amalgamations of social and technical elements that collectively contribute to goal-driven behavior, exerting influence on both social and technical stakeholders involved in the system Cooper and Foster (1971); Narayanan Venkit (2023). NLP applications are now prominently employed in diverse domains ranging from health and medicine Lee (2018) to policy-making Jin and Mihalcea (2022) and entertainment Werning (2024), thereby solidifying their pivotal role in everyday life.
@@
The proliferation of language models has also drawn substantial attention to its limitations and potential risks like the propagation of misinformation and the exacerbation of biases Bender et al. (2021); Gupta et al. (2024). Concurrently, there has been a surge in research focusing on examining the phenomenon of ‘hallucinations’ in language models Ji et al. (2023a). This trend is evident in the recent rise in peer-reviewed publications investigating hallucinations in NLP and language models, as depicted in Fig 1, sourced from SCOPUS.
In the domain of NLP, numerous frameworks have been proposed to conceptualize hallucination, with a predominant focus on its negative connotation. The negative notion commonly revolves around a model’s tendency to exhibit specific errors, particularly evident in tasks such as image captioning and text generation. Hallucination, in this context, denotes instances where the model generates references to non-existent objects or statements, despite lacking corresponding examples in the training dataset Ji et al. (2023a). Despite the increase in studies on hallucination, few works have underscored the need for a cohesive framework and precise definitions within NLP and language model research Filippova (2020a).
Acknowledging ‘hallucination’ as a social construct is imperative. Given the evolution of these models into sociotechnical systems, a pressing need arises to grasp the interdisciplinary nuances surrounding this phenomenon to understand how the field addresses it. This necessity is further underscored by research illustrating the societal repercussions of hallucinations on society Dahl et al. (2024). Therefore we see an increase in the need to understand how NLP defines and conceptualises hallucination. Based on this requirement, the following questions guide this study:
  • RQ1: What are the definitions and common frameworks used to explain hallucinations in NLG-published articles?
@@
To answer the RQ1, we first conduct an audit of the field of hallucinations in NLP by surveying 103 peer-reviewed articles111github.com/PranavNV/The-Thing-Called-Hallucination. Subsequently, we conduct a practitioner survey involving 171 researchers and academics within the field to explore their viewpoints on this phenomenon and answer RQ2. Through this work, we aim to understand the challenges posed by hallucinations in NLP. Finally, based on our findings, we construct an ethics framework to guide future endeavors to understand and address hallucinations in language models.
@@ 2 Evolution of Hallucination in NLP
In the modern deep learning era, hallucination was used first by Andrej Karpathy in his blog focusing on Recurrent Neural Networks Karpathy (2015). He used the term within the context of language models by illustrating how an LSTM could generate non-existent URLs, effectively ‘hallucinating’ data. The term then gained major traction with the launch of ChatGPT Wu et al. (2023), where it referred to inaccuracies and factual mistakes produced by models Ji et al. (2023a). However, the field lacked a unified definition, leading to a spectrum of interpretations Filippova (2020b). In one of the earlier works, Maynez et al. (2020) divides term usage into intrinsic and extrinsic hallucination. Intrinsic hallucinations are consequences of synthesizing content using the information present in the input. Extrinsic hallucinations are model generations that ignore the source material altogether.
Initial attempts at detecting hallucination in LMs try to calibrate model probabilities to tell itself if it is hallucinating Tian et al. (2023); Lin et al. (2022); Kadavath et al. (2022). Some attempts have been made to use transferable adversarial attacks to detect hallucination Yu et al. (2023).
Initial works on reducing hallucination aimed to retrain LMs with extra knowledge Zhang et al. (2019); Sun et al. (2021). Recently, self-checking, wherein models check their outputs for factual correctness, has shown progress as well Manakul et al. (2023); Luo et al. (2023). Another direction to reduce hallucination in LMs is by using external knowledge for producing outputs, commonly known as Retrieval Augmented Generation Systems (RAGs) Lewis et al. (2020); Guu et al. (2020); Zheng et al. (2023) whereby the model reduces hallucination using a controlled setting, such as querying on internal data.
@@
3 Social Perspectives on Hallucinations
Although the exploration of hallucination in NLP is solely technocentric, its conceptual roots and applications are deeply intertwined with societal interpretations. To gain a comprehensive understanding of the term ‘hallucination,’ it is crucial to consider its broader usage and implications beyond NLP.
Hallucination has been studied across disciplines like psychology and neurology Steele (2017); Macpherson and Platchias (2013); Legault (2020). Essentially, hallucinations involve “perceptions arising in the absence of any external reality – seeing or hearing things that are not there” Steele (2017). These experiences can be termed ‘veridical hallucinations,’ where “one hallucinates, but one’s experience is accurate” Macpherson and Platchias (2013). Although this definition is commonly used in NLP, often with negative connotations, hallucinations have a wide scope, originating from various fields such as psychology, neurology, and philosophy.
Hallucination and Medicine: Hallucination is believed to have neurological origins, often emerging from induced states such as drug usage, psychosis, sensory deprivation, or migraines Legault (2020). These experiences can encompass various sensory modalities like auditory, visual, olfactory, tactile, gustatory, or somatic sensations Boeving (2020). Modern neurological research indicates a correlation between the modality of hallucinations and specific brain regions responsible for processing associated sensory information Legault (2020). This suggests that while hallucinations may not align with external reality, they are linked to brain regions responsible for processing perceptions from the external world.
Hallucination and Creativity: Hallucinations are distinct from imagination, rumination, or dreaming. Some studies suggest that individuals prone to non-severe forms of hallucinations may exhibit heightened generative creativity Fink et al. (2014); Mason et al. (2021). However, it is essential to advocate for increased empirical rigor to establish strong connections between specific mental states that may lead to hallucinations and the creative thinking process Fink et al. (2014). Another commonly held belief revolves around using hallucinations as a means to access intuition, creativity, and innovative ways of thinking. This belief often extends to the use of external substances like psychedelics, which are believed to facilitate altered states of consciousness conducive to creative exploration Mason et al. (2021).
The exploration of various perceptions surrounding hallucination serves to explain its multifaceted interpretations. Contrary to prevailing negative assumptions in NLP, hallucination encompasses a spectrum of connotations. Utilizing the term ‘hallucination’ in the realm of NLP can lead to misconceptions for several reasons. Primarily, AI lacks sensory understanding, and the ‘errors’ in AI systems result from discrepancies in input data and prompts rather than a deficit in external stimuli. Secondly, this metaphor bears the potential to engender stigma by associating negative AI phenomena with specific aspects of mental illness Pal et al. (2023). Such linkage risks undermining endeavors aimed at destigmatizing psychiatric and mental health domains Maleki et al. (2024). Lastly, with machine learning models widely used in fields like medicine Ji et al. (2023c), a lack of contextual understanding of ‘hallucination’ may lead to misinterpretations of common terminology.
4 Critical Analysis of Hallucination in NLP
Given the discourse outlined in the preceding sections, it becomes evident that the framework employed for hallucination in AI substantially deviates from established social frameworks. This discrepancy presents challenges, as mentioned previously, when examining hallucination within sociotechnical systems Venkit et al. (2023), which strive to devise interdisciplinary solutions for a nuanced understanding of and for society. Therefore, conducting a critical analysis of this phenomenon becomes necessary to understand how the field of NLP conceptualizes and measures hallucination.
@@ 4 Critical Analysis of Hallucination in NLP
To accomplish this, we conducted an audit of works from the ACL anthology using specific keywords such as ‘hallucination’, ‘NLP (OR) AI’ AND ‘hallucinations’, ‘fabrication’, and ‘confabulations’. We surveyed papers released on and before March 19th, 2024. From this search, a total of 164 papers were retrieved. After filtering out papers that were not directly related to hallucination research or those that merely mentioned the term without substantial focus on the topic, we arrived at a corpus of 103 papers. This corpus forms the basis for our audit and analysis of hallucination research, specifically within the NLP domain.
@@ 4.1 Conceptualization of Hallucination
We performed an iterative thematic analysis Vaismoradi et al. (2013) to uncover the various applications of hallucination research in NLP. To ensure accuracy and prevent misclassification, this recursive process was employed. This resulted in the identification of seven distinct fields that address research on hallucination (as shown in Table 1).
Application Definitions Frequency
Conversational AI
AI systems designed for natural language conversations,
understanding inputs, and generating appropriate responses
38
Abstractive Summarization
Generating concise summaries by preserving main ideas
and context, often creating new sentence
16
Data2Text Generation
Automatically converting structured data into human-
readable text, used in reporting and narratives
14
Machine Translation
Automatically translating text between languages using
computational methods like neural networks
12
Image and Video Captioning
Generating descriptive captions for visual content, aiding
accessibility and understanding
8
Data Augmentation
Techniques to increase data diversity and quality, improving
model performance individual aspects of an entity
8
Miscellaneous
Encompasses additional non-accomodated tasks like natural
language inference and factuality detection
7
Table 1: Frameworks of Sentiment and corresponding definitions in Sentiment Analysis
This taxonomy affords insights into the pervasive nature of hallucination within this domain of literature. Notably, it reveals that hallucination in NLP transcends beyond text generation, extending its conceptualization to encompass broader domains such as Image-Video Captioning, Data Augmentation, and Data-to-Text Generation tasks. This underscores the significance of hallucination both within and beyond the realm of NLG. Moreover, our classification framework provides us with a faceted analysis of how each of these tasks defines the concept of hallucination.
@@
Large Language Models often exhibit a tendency to produce exceedingly confident, yet erroneous, assertions commonly referred to as hallucinations. - Zhang et al. (2023a)
@@
Hence, within the domain of NLP, a notable deficiency persists in grasping coherent characteristics of hallucination. This shortfall underscores the risk of potential misappropriation of the term when employed in divergent contexts.
@@
Table 2: The attributes that appear in the definitions of hallucination in the papers analyzed in the survey.
We now analyze what aspects of the definitions of hallucination most commonly occur within each of our identified sub-fields of NLP [^2] The breakdown of all the works associated with each of the subfields is in our Appendix. (Table 1).
Conversational AI: In this sub-field, hallucination encompasses fluency, non-factuality, and both intrinsic and extrinsic hallucinations. The definitions’ facets highlight that dialogue systems must balance conversational fluency with factual consistency, aligning both with prior conversation and real-world truths.
Abstractive Summarization: Works in this sub-field mainly focuses on extrinsic and intrinsic hallucinations in defining it. Some definitions also mention the faithfulness of the generation. Despite the challenges of aligning with real-world facts and source consistency, prioritizing alignment and adherence to the original material has been shown to be essential in these works.
Data2Text Generation: Hallucinations are classified into extrinsic and intrinsic types, similar to abstractive summarization. Here, alignment with the underlying data is emphasized as the more critical factor when compared to the language used in generating the text.
Machine Translation: Definitions of hallucination predominantly concentrate on extrinsic hallucination, with rare mentions of intrinsic hallucinations. This observation suggests a lesser concern for stylistic nuances in text generation within this field, with a greater emphasis on comprehending and conveying translated content accurately.
Image and Video Captioning: Models are expected to maintain consistency with the source while also incorporating real-world knowledge to address gaps and apply common sense. Consequently, the definition of hallucination in this context encompasses intrinsic, extrinsic, and non-factual elements, highlighting these requirements.
Data Augmentation:: Works from this domain often omit explicit definitions of hallucination, indicating a divergence in emphasis or a nascent exploration of this construct within this sub-field.
Miscellaneous: Encompassing tasks such as language inference and factuality detection, this category’s definitions of hallucination encompass aspects like factuality, intrinsic and extrinsic hallucination, fidelity, and nonsensicality. It’s evident that within these subfields, hallucination addresses both the stylistic aspects of model output and the fidelity and accuracy of generated content.
From the analysis of different subfields, it is evident that each perceives hallucination differently, emphasizing specific attributes such as factuality, fidelity, or linguistic styles like confidence, while potentially overlooking others. This diversity indicates that hallucination as a concept is still in its early stages in the field, with various frameworks emerging and a general lack of consensus regarding its definition and application. Furthermore, the lack of social aspects in hallucination discussions in these subfields contrasts with the broader understanding and research in fields like healthcare.
@@ 4.2 Audit of Frameworks
The preceding section delved into examining hallucination within the various domains of NLP, exposing a need for coherence in the prevailing approaches within the field. Building upon this revelation, our analysis now pivots to underscore the pressing need for a unified understanding of hallucination, particularly in light of the pervasive use of language models in both public discourse and interdisciplinary research endeavors.
We now scrutinize existing literature through a critical lens to unravel the dominant frameworks employed in defining hallucination while also assessing the extent to which these models accurately capture the essence of this phenomenon. This examination facilitates the outlook on transparency in the approach to hallucination in NLP, as well as the potential challenges it poses.
We start by looking at how many of the selected works explicitly define hallucination. Out of the 103 papers under consideration, just 44 (42.7%) provide a definition of the term, leaving the majority—59 papers or 57.3%—either altogether omitting their understanding of hallucination in the context of their research or providing no definition of a framework for the same. This lack of transparency is not only concerning but also underscores the need for clarity, especially given the varied interpretations of hallucination across different research domains.
Taking our scrutiny a step further, we investigate whether the works defining hallucination reference and acknowledge preexisting frameworks. It emerges that only 29 papers or 27% of the selected works explicitly acknowledge and adhere to established frameworks of hallucination, while the remainder 73% either loosely define the term or devise new definitions tailored to their specific research scope. This trend within the field underscores a lack of consensus on the conceptualization of hallucination, leading to disparate interpretations and a shortage of discourse on the subject.
We also audit the sociotechnical nature of the definitions of hallucination in NLP. Hallucination, as elucidated in Section 3, is inherently imbued with social dimensions, manifesting varied perspectives across different social contexts. Recognizing this complexity, it becomes important to leverage and integrate these social nuances into the shared lexicon of hallucination within machine learning and NLP. Moreover, given the evolution of language models into sociotechnical systems, adopting a sociotechnical approach becomes necessary, given that the term ‘hallucination’ is inherently a shared vocabulary within these domains.
Unfortunately, out of the 103 works examined, only 3 acknowledge the sociotechnical nature of hallucination, with none utilizing this framework to inform their approach. This oversight underscores a need for research to explore the sociotechnical dimensions inherent in hallucination, showcasing the current limited depth of understanding within the ML and AI communities.
@@ 4.3 Audit of Metrics
In the analysis of the 103 papers, we observed that 87 of these works dedicate efforts to measuring ‘hallucination’. This observation depicts the prevailing trend within NLP, emphasizing the significance of quantifying the concept of hallucination across diverse research endeavors. Building upon prior studies such as Ji et al. (2023a), our analysis categorizes the common approaches in NLP for quantifying hallucination into four major themes: Human Evaluation, Data-driven Metrics, Statistical Metrics, and Mixed Methodologies.
@@
Statistical Metrics approach entails the analysis of lexical features to assess information overlap and contradictions between generated and reference texts Ji et al. (2023a). This method calculates a hallucination score based on the degree of mismatch, with higher discrepancies indicating lower accuracy, factuality or faithfulness and hence, higher hallucination. Statistical scores such as BLUE, ROUGE, and Error Rate metrics are commonly used in this approach. Our findings reveal that 35.2% of the works that quantify hallucination opt for statistical metrics, employing 25 distinct metrics (e.g., BLEU, BERTScore, ROUGE1/2, F1, Perplexity, Cosine Similarity) developed for this purpose. This variability in measurement underscores the lack of a standardized approach within the research community for quantifying hallucination using statistical methods.
@@
This metrics audit reveals significant knowledge gaps and challenges across various approaches. Notably, established research highlights areas for improvement in standard methods for measuring hallucination. For instance, methodologies like CHAIR and metrics such as ROUGE scores exhibit instability in measuring hallucination due to the need for complex human-crafted parsing rules for exact matching, rendering them susceptible to errors Li et al. (2023). Criticisms also extend to human evaluation methods, which are prone to inaccuracies in gauging hallucination within these models Smith et al. (2022).
Beyond methodological criticisms, our survey uncovers a trend of employing numerous distinct metrics and approaches within these frameworks to categorize hallucinations. Over time, this has led to a diverse set of parameters for measuring hallucination, with a general lack of consensus on a standardized measurement approach. This issue is exacerbated by the multitude of diverse approaches to perceiving hallucination in NLP, further highlighting the absence of a unified method to address this challenge, especially as these models have now shifted to become a sociotechnical solution.

### 14 Sep 2024 v2

# An Audit on the Perspectives and Challenges of Hallucinations in NLP

@@ Abstract
We audit how hallucination in large language models (LLMs) is characterized in peer-reviewed literature, using a critical examination of 103 publications across NLP research. Through the examination of the literature, we identify a lack of agreement with the term ‘hallucination’ in the field of NLP. Additionally, to compliment our audit, we conduct a survey with 171 practitioners from the field of NLP and AI to capture varying perspectives on hallucination. Our analysis calls for the necessity of explicit definitions and frameworks outlining hallucination within NLP, highlighting potential challenges, and our survey inputs provide a thematic understanding of the influence and ramifications of hallucination in society.
@@ 1 Introduction
Recent advancements in Natural Language Processing (NLP) have expanded beyond traditional Machine Learning tools, evolving into sociotechnical systems that combine social and technical aspects to achieve specific goals Gautam et al. (2024); Narayanan Venkit (2023). They have now become integral in various domains such as health, policy-making, and entertainment, Jin and Mihalcea (2022); Werning (2024) showcasing their significant impact on daily life. However, language models (LM) exhibit negative behaviours such as hallucination and biases Bender et al. (2021); Gupta et al. (2024). This has catalyzed a surge in research investigating the phenomenon of hallucinations in NLP Ji et al. (2023a), reflected in the escalating number of peer-reviewed publications on the topic, as illustrated in Fig 1, sourced from SCOPUS.
@@
Within the NLP domain, various frameworks have emerged to define hallucination, primarily emphasizing its negative aspect. Hallucination here refers to the model’s production of references to non-existent objects or statements, lacking supporting examples in the training data Ji et al. (2023a). Despite the growing research on this topic, there is still a notable divide in our understanding, a lack of a unified framework, and a need for precise definitions Filippova (2020a).
The necessity to understand this gap is accentuated by research demonstrating the societal impacts of hallucinations Dahl et al. (2024). Hence, there is a growing need to explore how the field of NLP conceptualizes hallucination. In line with this imperative, the following questions guide this study:
  • RQ1: What are the definitions and frameworks used to explain hallucinations in NLP?
@@
To answer the RQ1, we first conduct an audit of the field of hallucinations in NLP by surveying 103 peer-reviewed articles [^1] https://github.com/PranavNV/The-Thing-Called-Hallucination. Subsequently, we conduct a survey to 171 researchers and academics in the field to gather their perspectives on this phenomenon, providing a novel contribution to the literature, addressing RQ2. By surveying NLP practitioners, the paper incorporates real-world perspectives, enriching the theoretical discussions with practical insights. This audit therefore aims to broaden the communities perspective by presenting practical insights from researchers employing these methods in their work. We also propose an ethical framework to guide future efforts in comprehending and mitigating hallucinations in LLMs.
@@ 2 Evolution of Hallucination in NLP
In the modern deep learning era, hallucination was used first by Andrej Karpathy in his blog focusing on Recurrent Neural Networks Karpathy (2015). He used the term within the context of LM by illustrating how an LSTM could generate non-existent URLs, effectively ‘hallucinating’ data. The term then gained major traction with the launch of ChatGPT Wu et al. (2023), where it referred to inaccuracies and factual mistakes produced by models Ji et al. (2023a). However, the field lacked a unified definition, leading to a spectrum of interpretations Filippova (2020b). In one of the earlier works, Maynez et al. (2020) divides term usage into intrinsic and extrinsic hallucination. Intrinsic hallucinations are consequences of synthesizing content using the information present in the input. Extrinsic hallucinations are model generations that ignore the source material altogether.
@@
3 Related Surveys on Hallucination
We now provide an overview of several key surveys in the realm of NLP focusing on the topic of hallucination and why our work addresses a relevant gap in the field. Starting with Ji et al. (2023a), this survey extensively delves into the advancements and challenges concerning hallucination in NLG, distinguishing between intrinsic and extrinsic frameworks of hallucination. Additionally, it sheds light on fundamental terms such as hallucination, faithfulness, and factuality, along with prevalent metrics for quantifying these phenomena. Rawte et al. (2023b) categorizes existing works within the domain of LMs, covering various aspects including methods for detecting hallucination, mitigation techniques, datasets used, and evaluation metrics. Zhang et al. (2023c) addresses the challenges of hallucination in LLMs by categorizing hallucinations into input-conflicting, context-conflicting, and fact-conflicting types, diverging from traditional viewpoints.
While these surveys offer insights into the current state of hallucination research, they do not pay attention to critical examinations of the field’s weaknesses arising from a lack of discourse in defining hallucination and challenges due to the same. This deficiency in discussion reflects the broader trend within the entire field. Therefore, our audit answers this gap by critically examining how we conceptualize hallucination. We aim to highlight the challenges stemming from these definitions and to further conduct a practitioner survey within the community to understand researchers’ and developers’ perspectives on this issue.
4 Critical Analysis of Hallucination in NLP Literature
This section is dedicated to conducting an audit of hallucination research within NLP, aiming to uncover its applications and subsequently identify the strengths and weaknesses in current literature.
@@ 4 Critical Analysis of Hallucination in NLP Literature
To accomplish this, we conducted an audit of works from the ACL anthology using specific keywords such as ‘hallucination’, ‘NLP (OR) AI’ AND ‘hallucinations’, ‘fabrication’, and ‘confabulations’. We surveyed papers released on and before April 19th, 2024. From this search, a total of 164 papers were retrieved. After filtering out papers that were not directly related to hallucination research or those that merely mentioned the term without substantial focus on the topic, we arrived at a corpus of 103 papers. This corpus forms the basis for our audit and analysis of hallucination research, specifically within the NLP domain.
@@ 4.1 Conceptualization of Hallucination
We performed an iterative thematic analysis Vaismoradi et al. (2013) to uncover the various applications of hallucination research in NLP. To ensure accuracy and prevent misclassification, this recursive process was employed. This resulted in the identification of seven distinct fields that address research on hallucination (as shown in Table. 1).
NLP Tasks Frequency
Conversational AI 38
Abstractive Summarization 16
Data-to-Text Generation 14
Machine Translation 12
Image-Video Captioning 8
Data Augmentation 8
Miscellaneous 7
Table 1: Frequency of papers reviewed for each thematically grouped NLP tasks.
This taxonomy affords insights into the pervasive nature of hallucination in NLP. Notably, it reveals that hallucination transcends beyond text generation, extending its conceptualization to encompass broader domains such as Image-Video Captioning, Data Augmentation, and Data-to-Text Generation tasks. This depicts the significance of hallucination both within and beyond the realm of NLP. Moreover, our classification framework provides us with a faceted analysis of how each of these tasks defines the concept of hallucination.
@@
LLMs often exhibit a tendency to produce exceedingly confident, yet erroneous, assertions commonly referred to as hallucinations. - Zhang et al. (2023a)
@@
Hence, within NLP, a notable deficiency persists in grasping coherent characteristics of hallucination. This shortfall underscores the risk of potential misappropriation of the term when employed in divergent contexts. An extensive analysis of hallucination for each of the mentioned NLP tasks and its definition is illustrated in the Appendix 10.2.
@@
Table 2: The attributes that appear in the definitions of hallucination.
@@ 4.2 Audit of Frameworks
We now scrutinize the dominant frameworks employed in defining hallucination while also assessing the extent to which these models accurately capture the phenomenon. We start by looking at how many of the selected works explicitly define hallucination. Out of the 103 papers, just 44 (42.7%) provide a definition of the term, leaving the majority—59 papers or 57.3%—either altogether omitting their understanding of hallucination in the context of their research or providing no definition or a framework. This lack of transparency is not only concerning but also underscores the need for clarity, especially given the varied interpretations of hallucination across different research domains.
Taking our scrutiny a step further, we investigate whether the works defining hallucination reference and acknowledge preexisting frameworks. It emerges that only 29 papers or 27% of the selected works explicitly acknowledge and adhere to established frameworks of hallucination, while the remainder 73% either loosely define the term or devise new definitions tailored to their specific research scope. This trend within the field shows a lack of consensus on the conceptualization of hallucination, leading to disparate interpretations and a shortage of discourse on the subject.
We also audit the sociotechnical nature of the definitions of hallucination in NLP. Hallucination (elucidated in Appendix 10.1) inherently contains social dimensions, creating varied perspectives across different social contexts. Moreover, given the evolution of LMs into social spaces, adopting a sociotechnical approach becomes necessary, given that the term ‘hallucination’ is inherently a shared vocabulary within these domains. Unfortunately, out of the 103 works examined, only 3 acknowledge the this nature of hallucination, with none utilizing this framework to inform their approach. This underscores a need for research to explore the sociotechnical dimensions inherent in hallucination, showcasing the limited depth of understanding within the ML and NLP communities.
@@ 4.3 Audit of Metrics
In the analysis of the 103 papers, we observed that 87 of these works dedicate efforts to measuring ‘hallucination.’ This observation depicts the prevailing trend within NLP, emphasizing the significance of quantifying the concept of hallucination across diverse research efforts. Building upon prior studies such as Ji et al. (2023a), our analysis categorizes the common approaches in NLP for quantifying hallucination into four major themes: Statistical Metrics, Data-driven Metrics, Human Evaluation, and Mixed Methodologies.
@@
Statistical metrics calculates a hallucination score based on the degree of mismatch, with higher discrepancies indicating lower accuracy, factuality or faithfulness and hence, higher hallucination Ji et al. (2023a). Statistical scores such as BLUE, ROUGE, and Error Rate metrics are commonly used in this approach. Our findings reveal that 35.2% of the works that quantify hallucination opt for statistical metrics, employing 25 distinct metrics (e.g., BERTScore, F1, Perplexity, Cosine Similarity) developed for this purpose. This variability underscores the lack of a standardized approach.
@@
The metrics audit reveals significant knowledge gaps and challenges across various approaches. Notably, established research highlights areas for improvement in standard methods for measuring hallucination. For instance, methodologies like CHAIR and metrics such as ROUGE scores exhibit instability in measuring hallucination due to the need for complex human-crafted parsing rules for exact matching, rendering them susceptible to errors Li et al. (2023). Criticisms also extend to human evaluation methods, which are prone to inaccuracies in gauging hallucination within these models Smith et al. (2022).
Beyond methodological criticisms, our audit uncovers a trend of employing numerous distinct metrics and approaches within these frameworks to categorize hallucinations. Over time, this has led to a diverse set of parameters for measuring hallucination, with a general lack of consensus on a standardized measurement approach. This issue further highlights the absence of a unified method, especially as these models have now shifted to become a sociotechnical solution Bender et al. (2021).